what is the greatest common factor of the monomials 18x2, 27x4, 30x3?
Found 2 solutions by rfer, benazir.sj@gmail.com:
Answer by rfer(16322)

3x^2(6)
3x^2(9x^2)
3x^2(10x)

Answer by benazir.sj@gmail.com(109)

18x^2 =3*3*2 *x^2
27x^4 =3*3*3*x^2 *x^2
30x^3 =3*5*2*x^2*x
GCF =3*x^2
=3x^2